OpenStudy (anonymous):

1.A right triangle has leg of b of length 7 and a hypotenuse of length 11. What is a length of the other leg, a? Round to the nearest tenth if necessary. (1 point) A. 2.0 B. 8.5 C. 13.0 D. 9.6 2. Find the length of a line segment CD with endpoint C at (-3,1) and endpoint D at (5,6). Round your answer to the nearest tenth, if necessary. (1 point) A 9.4 B 5.4 C 3.6 D 11.7 3.Find the midpoint of a segment FG with point F at (-6,4) and point G at (8,-2). (1 point) A (-7,3) B (7,-3) C (1,1) D (-1,-1)

OpenStudy (anonymous):

1. In a right triangle, (Hypotenuse)^2 = (Base)^2+(Perpendicular)^2 (11)^2 = (7)^2 +a^2 121 = 49 + a^2 a^2 = 121-49 a^2 = 72 a = 8.48 2. Length of line segment CD = sqrt (x2 - x1)^2 + (y2 - y1)^2 = sqrt (5+3)^2 + (6-1)^2 = sqrt (64+25) = sqrt 89 = 9.43 3. x- coordinate of midpoint= (x1+x2)/2 = (-6+8)/2 = 1 y-coordinate of midpoint= (y1+y2)/2 = (4-2)/2 = 1 Midpoint= (1,1) I hope that helps.. :)

OpenStudy (anonymous):

1.[7^{2} + x^{2} = 11^{2} 49 + x^{2} = 121 x^{2} = 72 x = 8.5 by the way [11^{2} is the hypotenuse

OpenStudy (anonymous):

i'm only in 8th grade